# Thought-Path Contrastive Learning via Premise-Oriented Data Augmentation for Logical Reading Comprehension

### v2

@@ Abstract
Logical reading comprehension is a challenging task that entails grasping the underlying semantics of text and applying reasoning to deduce the correct answer. Prior researches have primarily focused on enhancing logical reasoning capabilities through Chain-of-Thought (CoT) or data augmentation. However, previous work constructing chain-of-thought rationales concentrates solely on analyzing correct options, neglecting the incorrect alternatives. Addtionally, earlier efforts on data augmentation by altering contexts rely on rule-based methods, which result in generated contexts that lack diversity and coherence. To address these issues, we propose a Premise-Oriented Data Augmentation (PODA) framework. This framework can generate CoT rationales including analyses for both correct and incorrect options, while constructing diverse and high-quality counterfactual contexts from incorrect candidate options. We integrate summarizing premises and identifying premises for each option into rationales. Subsequently, we employ multi-step prompts with identified premises to construct counterfactual context. To facilitate the model’s capabilities to better differentiate the reasoning process associated with each option, we introduce a novel thought-path contrastive learning method that compares reasoning paths between the original and counterfactual samples. Experimental results on three representative LLMs demonstrate that our method can improve the baselines substantially across two challenging logical reasoning benchmarks (ReClor and LogiQA 2.0). The data and code are released at https://github.com/lalalamdbf/TPReasoner.
@@ 1 Introduction
Figure 1: Generating counterfactual context from incorrect candidate options.
Logical reasoning is a fundamental component of human cognition, essential for comprehending text and applying reasoning to deduce appropriate conclusions. Recently, challenging logical reasoning benchmarks have been proposed through machine reading comprehension (MRC) tasks (Yu et al. 2020; Liu et al. 2023a), which require models to derive the correct answer based on the given context, question and options. With the advent of large language models (LLMs), enhancing their capabilities in logical reasoning is a crucial step toward achieving strong artificial intelligence (Chollet 2019). Especially, the highly advanced model, GPT-4 (Achiam et al. 2023), has exhibited remarkable abilities to handle such tasks. However, a broad spectrum of open LLMs, including LLaMA2 (Touvron et al. 2023), Mistral (Jiang et al. 2023) and LLaMA3 (AI@Meta 2024), still fall short in logical reasoning, significantly trailing behind GPT-4. Consequently, improving the logical reasoning capabilities of community models has increasingly garnered the attention of many researchers (Liu et al. 2023b; Jiang et al. 2023).
Figure 2: A logical reasoning example from Reclor (Yu et al. 2020) dataset. The CoT rationale is annotated by GPT-3.5 or GPT-4. The content associated with the underlined phrases and the matching colored fonts is consistent. Due to space constraints, we refer to the specific reasoning process as {thought-path}.
@@
In view of above challenges, we propose a premises-oriented data augmentation (PODA) framework. As shown in Figures 1 and 2, the objective of PODA is to generate CoT rationales that include analyses for both correct and incorrect options, while also constructing counterfactual contexts based on incorrect candidate options. In Figure 2, analyses for both correct and incorrect options are presented in Analyze options. Besides, we incorporate summarizing premises and identifying premises for each option into rationales. Each option has a specific relationship with these premises—either supported, contradicted, or unrelated. PODA will create high-quality and diverse counterfactual contexts using multi-step prompts based on these premises and relationships. Furthermore, since supervised fine-tuning (SFT) focuses solely on individual instances, it lacks the comparison between different samples. For original and counterfactual samples, there are thought-paths that indicate similar and dissimilar reasoning processes associated with options. Therefore, we propose a thought-path contrastive learning approach, which specifically compares thought-paths across different samples, facilitating the model’s capabilities to better distinguish diverse reasoning paths. The main contributions of this paper are summarized as follows:
@@ 3.1 Premise-Oriented Data Augmentation
PODA initially creates analyses by forming thought-paths for both correct and incorrect options. Summarizing premises and identifying premises for each option are incorporated into CoT rationales, which are essential for generating new data. The core idea of it is to prompt a large language model through in-context learning to generate counterfactual data that can flip the current answer to a new answer (e.g., Option a $\rightarrow$ Option c). A context can be divided into Premises (the known information from the text), which have specific relationships with the options. The relationships are categorized into three types: supported, contradicted and unrelated. Utilizing the premise as a foundational unit, we can construct counterfactual samples based on these relationships.
@@
As illustrated in Figure 2, we design a structural CoT consisting of three steps: (1) Summarize Premises: Extract supporting statements from the context to serve as premises. (2) Analyze Options: Conduct a thorough evaluation of each option, elucidating the specific relationships between the options and premises, referred to as thought-path in our work. (3) Derive answer: Consolidate all thought-paths and determine the final answer. To guarantee a well-structured CoT, we utilize Few-shot-CoT for data collection. [^1] In some cases where the model generates an incorrect response, we provide the correct answer to assist in its self-correction and refinement of the reasoning process.
@@
Upon a combination of the initial three stages, we then implement correctness verification using Few-shot-CoT to filter out incorrect samples. The prompt and output format of this stage align with CoT Rationale Annotation. [^2] Summarizing premises and identifying premises for each option are optional, and we maintain these intermediate steps to ensure that the format of rationales for both original and counterfactual samples remains consistent. The potential mistakes of samples primarily stem from the following three aspects:
@@
By harnessing LLM’s self-assessment of its generated samples, PODA ensures that the counterfactual data is of high-quality and minimal noise.
@@ 5.3 Evaluation of Data Quality
A primary concern is whether the synthetic data accurately matches the correctly labeled option. In order to evaluate this, we choose five outstanding LLMs, including Mixtral-8$\times$7B-Instruct, GPT-3.5 (gpt-3.5-turo-0613), LLaMA2-70B-chat, LLaMA3-70B-Instruct and GPT-4 (gpt-4o). Non-GPT series models evaluate all counterfactual data. Due to budget constraints, a random subset of 200 samples from the generated dataset are evaluated by GPT-3.5 and GPT-4. We utilize 3-shot CoTs to evaluate the accuracy. As illustrated in Table 3, the accuracy assessed by LLaMA3-70B-Instruct and GPT-4 can reach 82.19% and 93% respectively, indicating PODA can generate high-quality counterfactual data. Moreover, accuracies for other models vary between approximately 75% and 79%, reflecting the significant challenges and complexities presented by these data. Overall, the generated data can have applicability in both training and evaluation domains.
@@
We compare PODA with a rule-based method, LReasoner (Wang et al. 2022), which constructs counterfactual contexts by modifying logical expressions according to logical laws. Two random subsets of 200 counterfactual samples were selected respectively. We ensure that the two subsets are derived from the same set of original samples for a fair comparison. These instances are evaluated by GPT-4 (gpt-4o) using four key metrics: Coherence (Is the context well-connected and logically consistent?), Clarity (Is the context clear and easy to understand?), Relevance (Does the context relate to the question and options?), and Diversity (How does the counterfactual context differ from the original one?). Each context was rated on a scale from 1 (poor) to 5 (excellent) for each metric. Our method attains average scores of 4.61 for Coherence, 4.36 for Clarity, 4.71 for Relevance, 3.18 for Diversity, and 4.22 Overall. In contrast, LReasoner achieves average scores of 2.98 for Coherence, 2.96 for Clarity, 4.63 for Relevance, 1.08 for Diversity, and 2.91 Overall. This comparison clearly demonstrates that the contexts generated by our method significantly outperform those produced by the rule-based method in both quality and diversity.
@@
We compare PODA with LogiCoT, which also utilizes GPT-4 for CoT rationales but focuses solely on analyzing the correct option. Similarly, two random subsets of CoT rationales were selected from the same contexts. These rationales are assessed by GPT-4 (gpt-4o) using four key metrics: Coherence (Is the CoT rationale logically consistent?), Completeness (Does it offer a thorough explanation for the reasoning?), Relevance (Does it directly and effectively addresses the context, question and options?), and Faithfulness (Is it factually correct and free from fabricated details?). Each rationale was rated on a scale from 1 (poor) to 5 (excellent) for each metric. PODA surpasses LogiCoT across four metrics, especially in Completeness. This suggests that incorporating the analysis of incorrect options into the rationales can enhance the quality of CoTs.

### v3

@@ Abstract
Logical reading comprehension is a challenging task that involves understanding the underlying semantics of text and applying reasoning to deduce the correct answer. Prior researches have primarily focused on enhancing logical reasoning capabilities through Chain-of-Thought (CoT) or data augmentation. However, previous work constructing chain-of-thought rationales concentrates solely on analyzing correct options, neglecting the incorrect alternatives. Addtionally, earlier efforts on data augmentation by altering contexts rely on rule-based methods, which result in generated contexts that lack diversity and coherence. To address these issues, we propose a Premise-Oriented Data Augmentation (PODA) framework. This framework can generate CoT rationales including analyses for both correct and incorrect options, while constructing diverse and high-quality counterfactual contexts from incorrect candidate options. We integrate summarizing premises and identifying premises for each option into rationales. Subsequently, we employ multi-step prompts with identified premises to construct counterfactual context. To facilitate the model’s capabilities to better differentiate the reasoning process associated with each option, we introduce a novel thought-path contrastive learning method that compares reasoning paths between the original and counterfactual samples. Experimental results on three representative LLMs demonstrate that our method can improve the baselines substantially across two challenging logical reasoning benchmarks (ReClor and LogiQA 2.0).
Codehttps://github.com/lalalamdbf/TPReasoner
@@ 1 Introduction
Figure 1: Generating counterfactual context from an incorrect candidate option.
Logical reasoning is a fundamental component of human cognition, essential for comprehending text and applying reasoning to deduce appropriate conclusions. Recently, challenging logical reasoning benchmarks have been proposed through machine reading comprehension (MRC) tasks (Yu et al. 2020; Liu et al. 2023a), which require models to derive the correct answer based on the given context, question and options. With the advent of large language models (LLMs), enhancing their capabilities in logical reasoning is a crucial step toward achieving strong artificial intelligence (Chollet 2019). Especially, the highly advanced model, GPT-4 (Achiam et al. 2023), has exhibited remarkable abilities to handle such tasks. However, a broad spectrum of open LLMs, including LLaMA2 (Touvron et al. 2023), Mistral (Jiang et al. 2023) and LLaMA3 (AI@Meta 2024), still fall short in logical reasoning, significantly trailing behind GPT-4. Consequently, improving the logical reasoning capabilities of community models has increasingly attracted the attention of many researchers (Liu et al. 2023b; Jiang et al. 2023).
Context: [Content of the context]
Question: [Content of the question]
Options: [Content of the options]
Summarize Premises:
1. [Premise 1]
2. [Premise 2]
3. [Premise 3]
Analyze Options:
(a) [Thought-path 1]
Identify Premises : Unrelated to the premises.
(b) [Thought-path 2]
Identify Premises: Supported by premises 2 and 3.
(c) [Thought-path 3]
Identify Premises: Unrelated to the premises.
(d) [Thought-path 4]
Identify Premises: Contradicted by premise 1.
[A summary of thought-paths]. Therefore, the optimal correct answer is (b).
Table 1: A logical reasoning example. The CoT rationale is annotated by GPT-3.5 or GPT-4. Due to space constraints, we refer to the specific reasoning process as [thought-path].
@@
In view of above challenges, we propose a premises-oriented data augmentation (PODA) framework. As shown in Figure 1 and Table 1, the objective of PODA is to generate CoT rationales that include analyses for both correct and incorrect options, while also constructing counterfactual contexts based on incorrect candidate options. In Table 1, analyses for both correct and incorrect options are presented in Analyze options. Besides, we incorporate summarizing premises and identifying premises for each option into rationales. Each option has a specific relationship with these premises—either supported, contradicted, or unrelated. PODA will create high-quality and diverse counterfactual contexts using multi-step prompts based on these premises and relationships. Furthermore, since supervised fine-tuning (SFT) focuses solely on individual instances, it lacks the comparison between different samples. For original and counterfactual samples, there are thought-paths that indicate similar and dissimilar reasoning processes associated with options. Therefore, we propose a thought-path contrastive learning approach, which specifically compares thought-paths across different samples, facilitating the model’s capabilities to better distinguish diverse reasoning paths. The main contributions of this paper are summarized as follows:
@@ 3.1 Premise-Oriented Data Augmentation
PODA initially creates analyses by forming thought-paths for both correct and incorrect options. Summarizing premises and identifying premises for each option are incorporated into CoT rationales, which are essential for generating new data. The core idea of it is to prompt a large language model through in-context learning to generate counterfactual data that can reverse the current answer to a new answer. A context can be divided into Premises (the known information from the text), which have specific relationships with the options. The relationships are categorized into three types: supported, contradicted and unrelated. Utilizing the premise as a foundational unit, we can construct counterfactual samples based on these relationships.
@@
As illustrated in Table 1, we design a structural CoT consisting of three steps: (1) Summarize Premises: Extract supporting statements from the context to serve as premises. (2) Analyze Options: Conduct a thorough evaluation of each option, clarifying the specific relationships between the options and premises, referred to as thought-path in our work. (3) Derive answer: Combine all thought-paths and determine the final answer. To guarantee a well-structured CoT, we utilize Few-shot-CoT for data collection. [^1] In some cases where the model generates an incorrect response, we provide the correct answer to assist in its self-correction and refinement of the reasoning process.
@@
Upon a combination of the initial three stages, we then implement correctness verification using Few-shot-CoT to filter out incorrect samples. The prompt and output format of this stage align with CoT Rationale Annotation. The potential mistakes of samples primarily stem from the following three aspects:
@@ 5.3 Evaluation of Data Quality
A primary concern is whether the synthetic data accurately matches the correctly labeled option. In order to evaluate this, we choose five outstanding LLMs, including Mixtral-8$\times$7B-Instruct, GPT-3.5 (gpt-3.5-turo-0613), LLaMA2-70B-chat, LLaMA3-70B-Instruct and GPT-4 (gpt-4o-2024-05-13). Non-GPT series models evaluate all counterfactual data. Due to budget constraints, a random subset of 200 samples from the generated dataset are evaluated by GPT-3.5 and GPT-4. We utilize 3-shot CoTs to evaluate the accuracy. As illustrated in Table 4, the accuracy assessed by LLaMA3-70B-Instruct and GPT-4 can reach 82.19% and 93% respectively, indicating PODA can generate high-quality counterfactual data. Moreover, accuracies for other models vary between approximately 75% and 79%, reflecting the significant challenges and complexities presented by these data. Overall, the generated data can have applicability in both training and evaluation domains.
@@
We compare PODA with a rule-based method, LReasoner (Wang et al. 2022), which constructs counterfactual contexts by modifying logical expressions according to logical laws. Two random subsets of 200 counterfactual samples were selected respectively. We ensure that the two subsets are derived from the same set of original samples for a fair comparison. These instances are evaluated by GPT-4 (gpt-4o-2024-05-13) using four key metrics: Coherence (Is the context well-connected and logically consistent?), Clarity (Is the context clear and easy to understand?), Relevance (Does the context relate to the question and options?), and Diversity (How does the counterfactual context differ from the original one?). Each context was rated on a scale from 1 (poor) to 5 (excellent) for each metric. Our method attains average scores of 4.61 for Coherence, 4.36 for Clarity, 4.71 for Relevance, 3.18 for Diversity, and 4.22 Overall. In contrast, LReasoner achieves average scores of 2.98 for Coherence, 2.96 for Clarity, 4.63 for Relevance, 1.08 for Diversity, and 2.91 Overall. This comparison clearly demonstrates that the contexts generated by our method significantly outperform those produced by the rule-based method in both quality and diversity.
@@
We compare PODA with LogiCoT, which also utilizes GPT-4 for CoT rationales but focuses solely on analyzing the correct option. Similarly, two random subsets of CoT rationales were selected from the same contexts. These rationales are assessed by GPT-4 (gpt-4o-2024-05-13) using four key metrics: Coherence (Is the CoT rationale logically consistent?), Completeness (Does it offer a thorough explanation for the reasoning?), Relevance (Does it directly and effectively addresses the context, question and options?), and Faithfulness (Is it factually correct and free from fabricated details?). Each rationale was rated on a scale from 1 (poor) to 5 (excellent) for each metric. PODA surpasses LogiCoT across four metrics, especially in Completeness. This suggests that incorporating the analysis of incorrect options into the rationales can enhance the quality of CoTs.
@@ 6 Conclusion
Acknowledgments
This work is supported by the grants from the National Natural Science Foundation of China (No. 62172044 and No. 62376130). The authors would like to thank the organizers of AAAI 2025 and the reviewers for their helpful suggestions.
@@ C.3 Prompts for Evaluating Data Quality
C.4 Human Evaluation
We supplement the human evaluation of data quality, with all evaluators being undergraduate or master’s students.
Method Coh Clar Rel Div Overall
LReasoner 2.44 3.17 4.37 1.07 2.76
PODA 4.06 4.39 4.44 3.09 4.00
Table 7: Human evaluation of counterfactual data on Coherence (Coh), Clarity (Clar), Relevance (Rel), and Diversity (Div).
Method Coh Comp Rel Faith Overall
LogiCoT 4.37 2.93 4.14 4.73 4.04
PODA 4.41 4.47 4.64 4.83 4.59
Table 8: Human evaluation of rationales on Coherence (Coh), Completeness (Comp), Relevance (Rel), and Faithfulness (Faith).